OpenStudy (anonymous):

A toy car having a mass of 4.00 x 10^-2 kilogram starts from rest at point A and travels 3.60 meters along a uniform track until coming to rest at Point B. Point A's height is 0.80 m. Point B's height is 0.50 m. Calculate the magnitude of the frictional force acting on the car. [Assume the frictional force is constant.]

OpenStudy (anonymous):

\[with conservation of energy..we can get: m.g.h=F.s (potential energy convert \to friction work)\] h=0.3 m s=3.6 m so, \[4\times10^{2} kg \times10 m/s^{2}\times0.3 m=F \times3.6 m\] then we get, F=1/30 N= 0.0334 N